In in a triangle ABC, angle B = pi/3
and Sin(A) Sin(C) = y
Fin the set of all possible values of y

Grade:11

1 Answers

Vijay Mukati
8 years ago
HInt. Since A + B + C = 180. Therfore, A+ B = 120 and A – C = 2(60 – C). Use the formula for 2SinASinC and substitute of above values. Thanks.